### v2

@@ 2.2 MedPromptX-VQA Dataset
The creation of MedPromptX-VQA involves three steps: (1) extraction of laboratory test results from the chartevents table within the MIMIC-IV dataset [8], resulting in 357 features in total, (2) feature engineering, which includes identification of the most strongly correlated features in relation to the label using Pearson method, and (3) transformation of clinical charts into textual representations using comma-separated values (i.e., serialization). Finally, the dataset is structured to support the in-context learning task, where each record has interleaved image and text, encompassing both positive and negative samples of patients. The motivation for feature selection is to maintain input consistency between the few-shot data and the query sample. This means that the features present in the query sample should already be represented by the candidates, while also adhering to the context length. Hence, we set a maximum of 10 features per label. The selected features are presented in Table 5 in Appendix.
@@ 3 Experimental Setup
We employed a randomized order strategy for the input sequences across several SOTA models; namely Med-Flamingo [15], OpenFlamingo [2], BioMedLM [3] and Clinical-T5-Large [14]. Moreover, the number of few-shot samples remained consistent at 6 across all the models and they were chosen randomly. For MedPromptX, the number of few-shot candidates is dynamically reduced by the DPS technique, resulting in a unique configuration for each query instance. The exclusion criteria for a candidate involve eliminating instances where the cosine similarity falls below a certain threshold, set at 70%. Furthermore, all experiments were conducted using NVIDIA A100-SXM GPU equipped with 40GB of dedicated memory. For MedPromptX, the frozen language encoder employed is LLaMA-7B [21], while the frozen visual encoder is CLIP ViT-L-14 [17]. Table 3 in Appendix shows detailed descriptions of the used models.
@@ 4.1 Ablations
Adjusting the threshold for DPS can significantly affect performance; an extremely high threshold restricts the model from including meaningful examples, while an extremely low threshold retains nearly the same examples (Figure 3 in Appendix). Moreover, the utilization of multimodal similarity (Table 4 in Appendix) enhances the instance selection process by capturing a more comprehensive representation of the data compared to single-modality approaches.
@@ Appendix 0.A Appendix
Table 3: Overview of large language models and visual-language models.
@@
Figure 3: Comparison of MedPromptX under different DPS thresholds.
Table 4: Comparison of employing DPS with averaged similarity scores from two modalities versus employing similarity scores based on a single modality.
@@
Table 5: Summary of the top correlated features that contribute to each label’s prediction, providing a clear understanding of the significant variables driving our model’s performance.

### v3

@@ 2.2 MedPromptX-VQA Dataset
The creation of MedPromptX-VQA involves three steps: (1) extraction of laboratory test results from the chartevents table within the MIMIC-IV dataset [8], resulting in 357 features in total, (2) feature engineering, which includes identification of the most strongly correlated features in relation to the label using Pearson method, and (3) transformation of clinical charts into textual representations using comma-separated values (i.e., serialization). Finally, the dataset is structured to support the in-context learning task, where each record has interleaved image and text, encompassing both positive and negative samples of patients. The motivation for feature selection is to maintain input consistency between the few-shot data and the query sample. This means that the features present in the query sample should already be represented by the candidates, while also adhering to the context length. Hence, we set a maximum of 10 features per label. The selected features are presented in Table 0.A.3 in Appendix.
@@ 3 Experimental Setup
We employed a randomized order strategy for the input sequences across several SOTA models; namely Med-Flamingo [15], OpenFlamingo [2], BioMedLM [3] and Clinical-T5-Large [14]. Moreover, the number of few-shot samples remained consistent at 6 across all the models and they were chosen randomly. For MedPromptX, the number of few-shot candidates is dynamically reduced by the DPS technique, resulting in a unique configuration for each query instance. The exclusion criteria for a candidate involve eliminating instances where the cosine similarity falls below a certain threshold, set at 70%. Furthermore, all experiments were conducted using NVIDIA A100-SXM GPU equipped with 40GB of dedicated memory. For MedPromptX, the frozen language encoder employed is LLaMA-7B [21], while the frozen visual encoder is CLIP ViT-L-14 [17]. Table 0.A.1 in Appendix shows detailed descriptions of the used models.
@@ 4.1 Ablations
Adjusting the threshold for DPS can significantly affect performance; an extremely high threshold restricts the model from including meaningful examples, while an extremely low threshold retains nearly the same examples (Figure 0.A.1 in Appendix). Moreover, the utilization of multimodal similarity (Table 0.A.2 in Appendix) enhances the instance selection process by capturing a more comprehensive representation of the data compared to single-modality approaches.
@@ Appendix 0.A Appendix
Table 0.A.1: Overview of large language models and visual-language models.
@@
Figure 0.A.1: Comparison of MedPromptX under different DPS thresholds.
Table 0.A.2: Comparison of employing DPS with averaged similarity scores from two modalities versus employing similarity scores based on a single modality.
@@
Table 0.A.3: Summary of the top correlated features that contribute to each label’s prediction, providing a clear understanding of the significant variables driving our model’s performance.